OpenStudy (anonymous):

cos(pi/2-x)=sin(x) True or False

OpenStudy (anonymous):

false, doesn't work out

OpenStudy (anonymous):

false

OpenStudy (zzr0ck3r):

\(\cos(A-B) = \cos(A)\cos(B)+\sin(A)\sin(B)\) So in fact \(\cos(\dfrac{\pi}{2}-x)=\cos(\dfrac{\pi}{2})\cos(x)+\sin(\dfrac{\pi}{2})\sin(x)=0*\cos(x)+1*\sin(x)=\sin(x)\)

OpenStudy (zzr0ck3r):

@plohrb @Dwayne_Johnson This is the problem with just giving answers...

OpenStudy (freckles):

this is also cofunction identity compare cos(pi-x) 's ratio and sin(x)'s ratio